Alexander runs his own canteen. This morning, he sold 13 "chef's plates" and 32 "quick and delicious" snacks. He made $143.25 on these sales. In the afternoon, he sold 7 "chef's plates" and 23 "quick and delicious" snacks, and collected $93. What are the costs of each meal?
Answer by mananth(16946)

x= cost of chef plates
y= cost of snacks
Total sold
13.00 x + 32.00 y = 143.25 .............1
Total value
7.00 x + 23.00 y = 93.00 .............2
Eliminate y
multiply (1)by -23.00
Multiply (2) by 32.00
-299.00 x -736.00 y = -3294.75
224.00 x + 736.00 y = 2976.00
Add the two equations
-75.00 x = -318.75
/ -75.00
x = 4.25
plug value of x in (1)
13.00 x + 32.00 y = 143.25
55.25 + 32.00 y = 143.25
32.00 y = 143.25 -55.25
32.00 y = 88.00
y = 2.75
x= 4.25 cost of chef plates
y= 2.75 cost of snacks
m.ananth@hotmail.ca